Consider three point charges at the corners of a triangle where q1 = 6.00 nC, q2 = -2.00 nC and q3 = 5.00 nC. Find the components of force F12, F32, resultant force on q2 in terms of magnitude and direction.
1
Expert's answer
2020-03-19T11:38:40-0400


As we are not provided with the length of the triangle so I have considered it as r

I have attached the diagram of the question

F1="\frac{kq_1q_2}{r^2}=\frac{9\times10^9\times6\times2\times10^{-18}}{r^2}=\frac{108}{r^2}nF"


F2="\frac{kq_1q_2}{r^2}=\frac{9\times10^9\times5\times2\times10^{-18}}{r^2}=\frac{90}{r^2}nF"


We will find the find the resultant of these two vectors using the formula

Fnet="\sqrt{F_1^2+F_2^2+2F_1F_2\cos60^o}"